solve for the Force of 2 charges, q1 = -63 nC and q2 = +42 nC with a distance of 5.5 cm. (Note: ϵ0 has a value of 8.85 x 10⁻¹² C2/ N m2) (C2/ N m2)


1
Expert's answer
2022-03-08T08:56:29-0500

The Coulomb's law says

"F=k\frac{q_1q_2}{r^2}"

"F=9*10^9*\frac{63*10^{-9}*42*10^{-9}}{(0.055)^2}=7.87*10^{-3}\:\rm N"